OpenStudy (anonymous):

The Louisiana Superdome covers an area of 363,000 square feet. Find the diameter of this round arena to the nearest foot.

OpenStudy (kropot72):

Area of a circle =\[\pi \times r ^{2}\] \[363000=\pi \times r ^{2}\] \[r=\sqrt{\frac{363000}{\pi}}\]

OpenStudy (kropot72):

Can you solve it now?

OpenStudy (anonymous):

no I'm sorry , i got 363000 div by 3.14 = 115605.....

OpenStudy (kropot72):

What is the square root of 115605?

OpenStudy (anonymous):

340 n change

OpenStudy (anonymous):

times it by 2

OpenStudy (kropot72):

That is the radius. Multiply by 2 to find the diameter.

OpenStudy (anonymous):

thank you